# Not All Samples Should Be Utilized Equally: Towards Understanding and Improving Dataset Distillation

### v1

@@ 3.3 Matching with Sample Difficulty Correction
~$\mathcal{L}_{\tilde{\mathcal{B}}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}^{c}}=\mathbb{E}_{(x,y)\in\tilde{\mathcal{B}}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}^{c}}\left[\mathcal{L}(x,y;\theta_{t})\right],\mathcal{L}_{\tilde{\mathcal{B}}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}^{c}}=\mathbb{E}_{(x,y)\in\tilde{\mathcal{B}}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}^{c}}\left[\mathcal{L}(x,y;\theta_{t})\right],$ (9)
@@
$\tilde{L}(\theta_{t})=\mathbf{D}\left(\nabla_{\theta}\mathcal{L}_{\tilde{\mathcal{B}}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}^{c}}\left(\theta_{t}\right),\nabla_{\theta}\mathcal{L}_{\tilde{\mathcal{B}}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}^{c}}\left(\theta_{t}\right)\right),$ (10)
@@
$L_{\lambda}(\theta_{t})=\underbrace{\mathbf{D}\left(\nabla_{\theta}\mathcal{L}_{\mathcal{B}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}^{c}}\left(\theta_{t}\right),\nabla_{\theta}\mathcal{L}_{\mathcal{B}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}^{c}}\left(\theta_{t}\right)\right)}_{\text{Gradient Matching Loss}}+\underbrace{\lambda\left\|\nabla_{\theta}\mathcal{L}_{\mathcal{B}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}^{c}}\right\|_{2}}_{\text{Gradient Norm Regularization}}$ (11)
@@
$L_{\lambda}(\theta^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}_{t})=\underbrace{\mathbf{D}\left(\theta_{t+M}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}},\theta_{t+N}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}\right)/{\mathbf{D}\left(\theta_{t+M}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}},\theta_{t}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}}\right)}}_{\text{Trajectory Matching Loss}}+\underbrace{\lambda\left\|\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}\right\|_{2}}_{\text{Gradient Norm Regularization}}$ (12)
By adding the gradient norm regularization, we can implicitly enforce current matching-based methods to mainly concentrate on synthesizing easy samples to achieve better synthetic data quality. We provide the detailed algorithm pseudocodes for GM- and TM-based methods in Appendix B.3.
@@ 4.1 Basic Settings
Datasets and baselines. For GM-based methods, we followed previous works to conduct experiments on MNIST [10], FashionMNIST [48], SVHN [36] datasets. We utilized current GM-based methods, including DC [52], DSA [51], and DSAC [25] as baselines. For TM-based methods, we followed the recent papers to use CIFAR-10, CIFAR-100 [21], and Tiny ImageNet [23] datasets. We performed experiments on current baselines including MTT [3], FTD [12], TESLA [8], and DATM [16]. We added our Sample Difficulty Correction (SDC) for all these baseline methods. To ensure a fair comparison, we employed identical hyperparameters for GM-based and TM-based methods with and without SDC while keeping all other variables constant, such as model architecture and augmentations. As per convention, for TM-based methods, we used max test accuracy, while for GM-based methods, we utilized the test accuracy from the last iteration. We also compared our methods with classical data pruning algorithms including Random, Herding [47], and Forgetting [45]. All hyperparameters are detailed in Appendix B.1.
Neural networks for distillation. Without specific mention, we used ConvNet as default to conduct experiments. Consistent with other previous methods, we used 3-layer ConvNet for CIFAR-10, CIFAR-100, MNIST, SVHN, and FashionMNIST, and 4-layer ConvNet for Tiny ImageNet.
@@ 4.2 Main Results
Generalization performance to other architectures. We evaluated the generalizability of synthetic datasets generated through distillation. Specifically, we used DSAC and DATM, which are current SOTA methods in GM-based and TM-based distillation, respectively. After distillation, the synthetic datasets were assessed using various neural networks, including ResNet-18 [17], VGG-11 [41], AlexNet [22], LeNet [24] and MLP. As shown in Table 3, even though our synthetic datasets were distilled using ConvNet, it generalizes well across most networks. Notably, for the experiment of DATM on CIFAR-10 with IPC = 1, employing SDC resulted in an accuracy improvement of 4.61% when using AlexNet. Employing SDC to DSAC led to an accuracy improvement of 0.9% on SVHN with IPC = 10 when using MLP. Additional results can be found in Appendix C.1.
@@ 4.3 Further Discussions
Discussion of SDC coefficient $\lambda$. The selection of the regularization coefficient $\lambda$ is pivotal for the quality of the distilled dataset. Our theory suggests that a larger $\lambda$ typically produces better synthetic datasets for smaller IPC values. Ideally, for low IPC settings, it is better to employ a large $\lambda$ to strongly penalize sample difficulty, whereas, for high IPC settings, the required $\lambda$ can be small or even close to zero in extreme cases. For simplicity and to maintain consistency across different datasets and baseline methods, we have set $\lambda=0.002$ as the default value in most of our experiments. As demonstrated in Figure 6, this choice of $\lambda$ aligns with the IPC values. Results for FTD and TESLA are based on CIFAR-10, results for DSA are based on SVHN, and results for DSAC are based on MNIST. Additionally, we further show that the choice of $\lambda$ is not sensitive in Appendix C.3.
Adaptive sample difficulty correction by adaptively increasing $\lambda$ during distillation. While our SDC seeks simplicity in regularization, DATM [16] claims that the matching difficulty is increased through optimization. Inspired by their observation, we implemented a strategy where $\lambda$ increases progressively throughout the matching phases. This method is designed to incrementally adjust the focus from easier to more complex patterns. Inspired by their observation, we applied an Adaptive Sample Difficulty Correction (ASDC) strategy in our experiments with a TM-based method on the CIFAR-100 with IPC = 1 and with a GM-based method on the FashionMNIST with IPC = 1. The $\lambda$ of DATM was initialized to 0.02 and logarithmically increased to 0.08 over 10,000 iterations and DSAC was initialized to 0.002 and logarithmically increased to 0.008 over 10,000 steps. For DATM, we use max test accuracy, while for DSAC, we use test accuracy. Experimental results of ASDC validate its potential to significantly enhance learning by finetuning regularization according to the complexity of the learned patterns. Figure 6 illustrates that ASDC further improves our method within SOTA matching methods. Additional results are provided in Appendix C.2.
@@ 5 Conclusion
Appendix A More Result on the Relationships between Sample Difficulty, Gardient Norm, and Loss
@@
Appendix B More Details of Experiments
B.1 Parameter Tables
B.1.1 GM-based Methods
@@ B.2 Limitation
B.3 Pseudocodes of adding SDC on Matching-based Distillation Methods
@@
Appendix C Exploring the Effectiveness of SDC in Additional Experiments
C.1 More Results on the Cross-architecture Evaluation
@@
C.2 More Results on the Adaptive Sample Difficulty Correction
@@
C.3 Sensitivity Analysis of SDC coefficient $\lambda$.
@@
Appendix D Analytical Theory for Dataset Distillation
@@
We provide saddle point equations for the cosine similarity $R$ between the probe $\theta^{\mathcal{D}_{\mathsf{probe}}}$ and the expert $\theta^{\mathcal{D}_{\mathsf{real}}}$, which will be discussed in Section D.1 and Section D.2. For our simulations, we set the parameter dimension $d=200$ for perfect probe settings, and set $d=50$ for imperfect probe settings. We averaged 100 simulation results to verify the theory.
D.1 Perfect Expert-Teacher Settings
The solution is given by the following saddle point equations for perfect exeprt-teacher settings, i.e., $\gamma=0$. For any given $\alpha_{\mathsf{syn}}$, these equations can be solved for the order parameters $R,\kappa$. From these parameters, the generalization error can be computed as $\varepsilon=\cos^{-1}(R)/\pi$.
$R=\frac{2\alpha_{\mathsf{syn}}}{f\sqrt{2\pi}\sqrt{1-R^{2}}}\int_{-\infty}^{\kappa}Dt\ \exp\bigg{(}-\frac{R^{2}t^{2}}{2(1-R^{2})}\bigg{)}\bigg{[}1-\exp\bigg{(}-\frac{\gamma(\gamma-2Rt)}{2(1-R^{2})}\bigg{)}\bigg{]}(\kappa-t)$ (13)
$1-R^{2}=\frac{2\alpha_{\mathsf{syn}}}{f}\int_{-\infty}^{\kappa}Dt\ \bigg{[}H\bigg{(}-\frac{Rt}{\sqrt{1-R^{2}}}\bigg{)}-H\bigg{(}-\frac{Rt-\gamma}{\sqrt{1-R^{2}}}\bigg{)}\bigg{]}(\kappa-t)^{2}$ (14)
@@ D.1 Perfect Expert-Teacher Settings
D.2 Imperfect Expert-Teacher Settings
We have shown the perfect student settings in Section D.1. When the probe student does not exactly match the expert, an additional parameter $\theta$ characterizes the angle between the probe student and the expert. Furthermore, an additional order parameter $\rho=\theta^{\mathcal{D}_{\mathsf{real}}^{\top}}\theta^{\mathcal{D}_{\mathsf{syn}}}$ represents the typical student-probe overlap, which must be optimized. Consequently, we derive three saddle point equations.
[…]
$\quad\quad\quad+\frac{1}{2\pi\Lambda}\exp\left(-\frac{\Delta(t,z)}{2\Lambda^{2}}\right)\bigg{(}\frac{\rho R-\cos\gamma}{1-\rho^{2}}\bigg{)}(\kappa-t)\bigg{>}_{z}$

### v2

@@ 3.3 Matching with Sample Difficulty Correction
~$\mathcal{L}_{\tilde{\mathcal{B}}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}^{c}}=\mathbb{E}_{(x,y)\in\tilde{\mathcal{B}}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}^{c}}\left[\mathcal{L}(x,y;\theta_{t})\right],$ (9)
~$\mathcal{L}_{\tilde{\mathcal{B}}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}^{c}}=\mathbb{E}_{(x,y)\in\tilde{\mathcal{B}}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}^{c}}\left[\mathcal{L}(x,y;\theta_{t})\right],$
@@
$\tilde{L}(\theta_{t})=\mathbf{D}\left(\nabla_{\theta}\mathcal{L}_{\tilde{\mathcal{B}}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}^{c}}\left(\theta_{t}\right),\nabla_{\theta}\mathcal{L}_{\tilde{\mathcal{B}}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}^{c}}\left(\theta_{t}\right)\right),$ (10)
@@
$L_{\lambda}(\theta_{t})$ $=\underbrace{\mathbf{D}\left(\nabla_{\theta}\mathcal{L}_{\mathcal{B}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}^{c}}\left(\theta_{t}\right),\nabla_{\theta}\mathcal{L}_{\mathcal{B}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}^{c}}\left(\theta_{t}\right)\right)}_{\text{Gradient Matching Loss}}$ (11)
$+\underbrace{\lambda\left\|\nabla_{\theta}\mathcal{L}_{\mathcal{B}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}^{c}}\right\|_{2}}_{\text{Gradient Norm Regularization}}$
@@
$L_{\lambda}(\theta^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}_{t})$ $=\underbrace{\mathbf{D}\left(\theta_{t+M}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}},\theta_{t+N}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}\right)/{\mathbf{D}\left(\theta_{t+M}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}},\theta_{t}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}}\right)}}_{\text{Trajectory Matching Loss}}$ (12)
$+\underbrace{\lambda\left\|\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}\right\|_{2}}_{\text{Gradient Norm Regularization}}$
By adding the gradient norm regularization, we can implicitly enforce current matching-based methods to mainly concentrate on synthesizing easy samples to achieve better synthetic data quality. We provide the algorithm pseudocodes for GM- and TM-based methods in Appendix 2.3.
@@ 4.1 Basic Settings
Datasets and baselines. For GM-based methods, we followed previous works to conduct experiments on MNIST [10], FashionMNIST [47], SVHN [33] datasets. We utilized current GM-based methods, including DC [52], DSA [51], and DSAC [23] as baselines. For TM-based methods, we followed the recent papers to use CIFAR-10, CIFAR-100 [19], and Tiny ImageNet [21] datasets. We performed experiments on current baselines including MTT [3], FTD [12], TESLA [8], and DATM [15]. We added our Sample Difficulty Correction (SDC) for all these baseline methods. To ensure a fair comparison, we employed identical hyperparameters for GM-based and TM-based methods with and without SDC while keeping all other variables constant, such as model architecture and augmentations. As per convention, for TM-based methods, we used max test accuracy, while for GM-based methods, we utilized the test accuracy from the last iteration. We also compared our methods with classical data pruning algorithms including Random, Herding [46], and Forgetting [41]. All hyperparameters are detailed in Appendix 2.1.
Neural networks for distillation.We used ConvNet as default to conduct experiments. Consistent with other previous methods, we used 3-layer ConvNet for CIFAR-10, CIFAR-100, MNIST, SVHN, and FashionMNIST, and 4-layer ConvNet for Tiny ImageNet.
@@ 4.2 Main Results
Generalization performance to other architectures. We evaluated the generalizability of synthetic datasets generated through distillation. We used DSAC and DATM, which are current SOTA methods in GM-based and TM-based distillation, respectively. After distillation, the synthetic datasets were assessed using various neural networks, including ResNet-18 [16], VGG-11 [37], AlexNet [20], LeNet [22] and MLP. As shown in Table 3, even though our synthetic datasets were distilled using ConvNet, it generalizes well across most networks. Notably, for the experiment of DATM on CIFAR-10 with IPC = 1, employing SDC resulted in an accuracy improvement of 4.61% when using AlexNet. Employing SDC to DSAC led to an accuracy improvement of 0.9% on SVHN with IPC = 10 when using MLP. Additional results can be found in Appendix 3.1.
@@ 4.3 Further Discussions
Discussion of SDC coefficient $\lambda$. The selection of the regularization coefficient $\lambda$ is pivotal for the quality of the distilled dataset. Our theory suggests that a larger $\lambda$ typically produces better synthetic datasets for smaller IPC values. Ideally, for low IPC settings, it is better to employ a large $\lambda$ to strongly penalize sample difficulty, whereas, for high IPC settings, the required $\lambda$ can be small or even close to zero in extreme cases. For simplicity and to maintain consistency across different datasets and baseline methods, we have set $\lambda=0.002$ as the default value in most of our experiments. As demonstrated in Figure 6, this choice of $\lambda$ aligns with the IPC values. Results for FTD and TESLA are based on CIFAR-10, results for DSA are based on SVHN, and results for DSAC are based on MNIST. Additionally, we further show that the choice of $\lambda$ is not sensitive in Appendix 3.3.
Adaptive sample difficulty correction by adaptively increasing $\lambda$ during distillation. While our SDC seeks simplicity in regularization, DATM [15] claims that the matching difficulty is increased through optimization. Inspired by their observation, we implemented a strategy where $\lambda$ increases progressively throughout the matching phases. This method is designed to incrementally adjust the focus from easier to more complex patterns. Inspired by their observation, we applied an Adaptive Sample Difficulty Correction (ASDC) strategy in our experiments with a TM-based method on the CIFAR-100 with IPC = 1 and with a GM-based method on the FashionMNIST with IPC = 1. The $\lambda$ of DATM was initialized to 0.02 and logarithmically increased to 0.08 over 10,000 iterations and DSAC was initialized to 0.002 and logarithmically increased to 0.008 over 10,000 steps. For DATM, we use max test accuracy, while for DSAC, we use test accuracy. Experimental results of ASDC validate its potential to significantly enhance learning by finetuning regularization according to the complexity of the learned patterns. Figure 6 illustrates that ASDC further improves our method within SOTA matching methods. Additional results are provided in Appendix 3.2.
@@ 5 Conclusion
Supplementary Material
1 More Result on the Relationships between Sample Difficulty, Gardient Norm, and Loss
@@
2 More Details of Experiments
2.1 Parameter Tables
2.1.1 GM-based Methods
@@ 2.2 Limitation
2.3 Pseudocodes of adding SDC on Matching-based Distillation Methods
@@
3 Exploring the Effectiveness of SDC in Additional Experiments
3.1 More Results on the Cross-architecture Evaluation
@@
3.2 More Results on the Adaptive Sample Difficulty Correction
@@
3.3 Sensitivity Analysis of SDC coefficient .
@@
4 Analytical Theory for Dataset Distillation
@@
We provide saddle point equations for the cosine similarity $R$ between the probe $\theta^{\mathcal{D}_{\mathsf{probe}}}$ and the expert $\theta^{\mathcal{D}_{\mathsf{real}}}$, which will be discussed in Section 4.1 and Section 4.2. For our simulations, we set the parameter dimension $d=200$ for perfect probe settings, and set $d=50$ for imperfect probe settings. We averaged 100 simulation results to verify the theory.
4.1 Perfect Expert-Teacher Settings
The solution is given by the following saddle point equations for perfect expert-teacher settings, i.e., $\gamma=0$. For any given $\alpha_{\mathsf{syn}}$, these equations can be solved for the order parameters $R,\kappa$. From these parameters, the generalization error can be computed as $\varepsilon=\cos^{-1}(R)/\pi$.
$R$ $=\frac{2\alpha_{\mathsf{syn}}}{f\sqrt{2\pi}\sqrt{1-R^{2}}}\int_{-\infty}^{\kappa}Dt\ \exp\left(-\frac{R^{2}t^{2}}{2(1-R^{2})}\right)$
×$\quad\times\left[1-\exp\left(-\frac{\gamma(\gamma-2Rt)}{2(1-R^{2})}\right)\right](\kappa-t)$
$1-R^{2}$ $=\frac{2\alpha_{\mathsf{syn}}}{f}\int_{-\infty}^{\kappa}Dt\ \bigg{[}H\left(-\frac{Rt}{\sqrt{1-R^{2}}}\right)$
$\quad-H\left(-\frac{Rt-\gamma}{\sqrt{1-R^{2}}}\right)\bigg{]}(\kappa-t)^{2}$
@@ 4.1 Perfect Expert-Teacher Settings
4.2 Imperfect Expert-Teacher Settings
We have shown the perfect student settings in Section 4.1. When the probe student does not exactly match the expert, an additional parameter $\theta$ characterizes the angle between the probe student and the expert. Furthermore, an additional order parameter $\rho=\theta^{\mathcal{D}_{\mathsf{real}}^{\top}}\theta^{\mathcal{D}_{\mathsf{syn}}}$ represents the typical student-probe overlap, which must be optimized. Consequently, we derive three saddle point equations.
$\frac{R-\rho\cos\gamma}{\sin^{2}\gamma}$ $=\frac{\alpha_{\mathsf{syn}}}{\pi\Lambda}\left\langle\int_{-\infty}^{\kappa}dt\ \exp\left(-\frac{\Delta(t,z)}{2\Lambda^{2}}\right)\right.$
×$\quad\left.\times(\kappa-t)\right\rangle_{z}$
$1-\frac{\rho^{2}+R^{2}-2\rho R\cos\gamma}{\sin^{2}\gamma}$ $=2\alpha_{\mathsf{syn}}\left\langle\int_{-\infty}^{\kappa}dt\ \frac{e^{-\frac{(t-\rho z)^{2}}{2(1-\rho^{2})}}}{\sqrt{2\pi}\sqrt{1-\rho^{2}}}\right.$
×$\quad\left.\times H\left(\frac{\Gamma(t,z)}{\sqrt{1-\rho^{2}}\Lambda}\right)(\kappa-t)^{2}\right\rangle_{z}$
$\frac{\rho-R\cos\gamma}{\sin^{2}\gamma}$ $=2\alpha_{\mathsf{syn}}\left\langle\int_{-\infty}^{\kappa}dt\ \frac{e^{-\frac{(t-\rho z)^{2}}{2(1-\rho^{2})}}}{\sqrt{2\pi}\sqrt{1-\rho^{2}}}\right.$
×$\quad\left.\times H\left(\frac{\Gamma(t,z)}{\sqrt{1-\rho^{2}}\Lambda}\right)\left(\frac{z-\rho t}{1-\rho^{2}}\right)(\kappa-t)\right\rangle_{z}$
$\quad+\frac{1}{2\pi\Lambda}\left\langle\exp\left(-\frac{\Delta(t,z)}{2\Lambda^{2}}\right)\right.$
×$\quad\left.\times\left(\frac{\rho R-\cos\gamma}{1-\rho^{2}}\right)(\kappa-t)\right\rangle_{z}$